350.0 mL of 0.450 mol/L sodium carbonate (Na2CO3) solution and 400.0 mL of 0.700 mol/L calcium chloride (CaCl2) solution are mixed together. Calculate the mass of calcium carbonate that would precipitate.


1
Expert's answer
2022-05-30T20:42:03-0400

Na2CO3 + CaCl2 = 2 NaCl + CaCO3

1. Na2CO 3va CaCl2 ning moddaning (mol) miqdorini topamiz. buning uchun quyidagi formuladan foydalanamiz:

n = CM * V

in the formula,

n — amount of substance (mol);

CM — concentration of molarity (mol/l);

V — volume of solution (litres).

a) Na2CO3 :

n = 0,45 mol/l * 0,35 l = 0,1575 mol

b) CaCl2 :

n = 0,7 mol/l * 0,4 l = 0,28 mol

2. To find the product of the reaction, we use to calculate the amount of substance (mol) that is less than the reactants. because a substance with a small amount (mol) is completely consumed in the reaction, and a substance with a high content (mol) increases by a certain amount. that is, we calculate using the amount of Na2CO3 and find the mass of CaCO3 :

According to above the reaction,

1mol Na2CO3 — 100gr (1mol) CaCO3

0,1575 mol Na2CO3 — X gr CaCO3

X = 0,1575 * 100 / 1 = 15,75 gr

ANSWER: 15,75gr CaCO3